a manufacturer estimates that D(p) = 3000e^-0.05p units of a particular good will be sold at market price of p cedis per unit. determine the market price that will result in marginal revenue of zero.


1
Expert's answer
2022-02-01T15:30:47-0500

The market price that will result in marginal revenue of zero is:

"MR = TR'(p) = (p\u00d7D(p))' = -150e^{-0.05p} p^2 + 3000e^{-0.05p} = 0,"

"e^{-0.05p}\u00d7(20 - p^2) = 0,"

"e^{-0.05p} = 0" (has no roots) or "p^2 = 20,"

p = 4.47.